For the function f(x) = x 2 use the Intermediate Value Theorem to prove that there exists a number c such that f(c) = 2. Your are proving the existence of the real number √ 2. Since f(1) < 2 < f(2) you can restrict f(x) to the interval [1, 2]


1
Expert's answer
2021-09-30T08:47:37-0400

"Here" "f(x)=x^2" .It is defined on R .Let us consider an interval [1,2]

Here "f(1)=(1)^2=1" , "f(2)=(2)^2=4"

According to intermediate value theorem there exists an element k satisfying f(a)<k<f(b) then f(c)=k where c belongs to (a , b) . Here 2 "\epsilon" (1,4) .So f(c)=2 "\implies" "c^2=2\implies" "c=\sqrt2" (proved)